### v2

# Can Open-source LLMs Enhance Data Synthesis for Toxic Detection?: An Experimental Study

Abstract
Effective toxic content detection relies heavily on high-quality and diverse data, which serves as the foundation for robust content moderation models. This study explores the potential of open-source LLMs for harmful data synthesis, utilizing prompt engineering and fine-tuning techniques to enhance data quality and diversity. In a two-stage evaluation, we first examine the capabilities of six open-source LLMs in generating harmful data across multiple datasets using prompt engineering. In the second stage, we fine-tune these models to improve data generation while addressing challenges such as hallucination, data duplication, and overfitting. Our findings reveal that Mistral excels in generating high-quality and diverse harmful data with minimal hallucination. Furthermore, fine-tuning enhances data quality, offering scalable and cost-effective solutions for augmenting datasets for specific toxic content detection tasks. These results emphasize the significance of data synthesis in building robust, standalone detection models and highlight the potential of open-source LLMs to advance smaller downstream content moderation systems. We implemented this approach in real-world industrial settings, demonstrating the feasibility and efficiency of fine-tuned open-source LLMs for harmful data synthesis.
Warning: This paper has instances of hateful and offensive language to serve as examples.
1 Introduction
Figure 1: Our experiment design is detailed: we first conducted prompt engineering on the models (stage 1), and then, for better results, we selected a subset of models for fine-tuning (stage 2).
In toxic content detection, (1) labeling real-world datasets is both expensive and labor-intensive, with the substantial volume of data required further compounding the challenge—our costs have reached millions of dollars. (2) Ensuring dataset diversity and quality remains a persistent issue, often hindering the development of robust detection models. (3) The inherent ambiguity and subjectivity of the term "harmful"—which varies across topics and annotator perspectives—adds another layer of complexity. Data synthesis presents a promising approach to mitigating these challenges. Moreover, many product applications demand real-time toxic content detection with stringent latency requirements (e.g., latency $\leq$ 0.01 seconds). Directly using SOTA Podolak et al. (2024); Fillies and Paschke (2024) LLM-based approaches might not reliably meet these constraints, making LLM-generated data synthesis for training smaller, more robust classification models a practical solution.
This study explores the use of prompt engineering and fine-tuning techniques with open-source LLMs to improve harmful data synthesis. Existing methods, such as ToxiGen Hartvigsen et al. (2022) and Toxicraft Hui et al. (2024) provide useful frameworks but rely on GPT-based frameworks for data generation, which face issues of cost and scalability. For example, ToxiGen employs GPT3 within the ALICE framework, producing high-quality datasets but at a high computational expense. These limitations highlight the need for alternative solutions. Our study addresses these challenges by investigating how open-source LLMs can facilitate harmful data synthesis. Open-source models offer comparable effectiveness to proprietary systems while reducing computational costs and enhancing explainability Ding et al. (2024). Many products can benefit from standalone data synthesis modules that generate training data for domain-specific and time-sensitive content safety applications, making open-source LLMs a scalable and cost-effective choice.
Our experimental approach includes a comprehensive evaluation of six open-source LLMs: Mistral Jiang et al. (2023), LLaMa2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024), explored across diverse prompt engineering strategies Schulhoff et al. (2024); Li (2023) and fine-tuning configurations. The research is divided into two key stages: In the first stage, we focus on prompt engineering by comparing LLMs, designing effective prompts, and experimenting with few-shot techniques to balance quality and diversity in the generated data Ding et al. (2024). However, we observed that relying solely on prompt engineering often results in suboptimal outputs due to the model’s inherent safety alignments, which limit harmful data generation. Consequently, we conducted a second stage of experiments involving fine-tuning. In the second stage, we fine-tuned the LLMs using proprietary datasets and explored various configurations, such as epoch settings and data mixing strategies. Our results demonstrate that fine-tuning effectively mitigates issues like hallucination in LLMs and enhances the quality of harmful data generation. However, challenges such as data duplication and overfitting were identified, emphasizing the need for a careful balance between training epochs and dataset diversity.
In this study, we do not aim to develop a new state-of-the-art model in toxic content detection but rather to explore and experiment with intuitive, effective ideas tailored for industrial applications of toxic content moderation systems. These efforts provide key insights into optimizing open-source LLMs for harmful data synthesis, enhancing both the robustness and accuracy of toxic content detection systems. Additionally, this study validates the practicality and efficiency of fine-tuned open-source LLMs in real-world industrial applications. Our main contributions are summarized as follows:
  • Our study is one of the first (if not the first) to apply prompt engineering and fine-tuning techniques on open-source LLMs, exploring their potential for harmful data synthesis and improving toxic content detection.
  • We show that fine-tuned open-source LLMs deliver cost-effective and scalable solutions for automated content moderation, providing actionable insights for developing larger and more robust harmful content detection systems through synthetic data generation.
  • We also offer valuable insights into the practicalities of production deployment, highlighting real-world challenges and solutions.
2 Methodology
Our approach utilizes prompt engineering and fine-tuning techniques on open-source LLMs to enhance harmful data synthesis, thereby improving toxic content detection systems. The methodology comprises two key stages: prompt engineering and fine-tuning, as illustrated in Figure 1.
2.1 Prompt Engineering
Given an open-source LLM $\mathcal{M}$, we design prompts $\mathcal{P}$ to generate harmful data $\mathcal{D}_{\text{harmful}}$. The prompt template is defined as:
[…]
where Role defines the context, Requirement specifies the task, and Few-Shot Examples provide examples to guide the model. In the initial phase of our study, we focused solely on prompt engineering. This involved creating various prompt templates tailored to elicit specific types of harmful data. Each prompt was carefully crafted following work by Zhou et al. (2023) to ensure clarity and relevance to the desired output. Despite the meticulous design of these prompts, we encountered significant challenges. The primary issue was the model’s inherent safety alignments, which are explicitly designed to prevent harmful content generation. While these alignments are crucial for ethical AI usage, they posed a hurdle for our specific goal of generating harmful data for augmentation purposes. Consequently, the generated data often lacked sufficient quality and diversity, hindering their utility in effective toxic content detection. To overcome these challenges, we experimented with various configurations, such as altering few-shot examples and adjusting requirement specificity. However, these adjustments yielded limited success in bypassing the internal safety mechanisms. After testing, the final prompt template are showed in Appendix A. Our findings showed that relying solely on prompt engineering was inadequate for our objectives. This realization prompted the integration of a second stage—fine-tuning—to enhance the model’s ability to generate high-quality harmful data by retraining its weights with carefully curated datasets.
2.2 Fine-Tuning
Based on the results from prompt engineering, we hypothesized that fine-tuning would yield a more effective solution. To achieve this, we employed LoRA methods Hu et al. (2022) for fine-tuning, updating the weights of $\mathcal{M}$ using the training dataset $\mathcal{D}_{\text{train}}$ to enhance the quality and diversity of $\mathcal{D}_{\text{harmful}}$. The fine-tuning objective, denoted as $\mathcal{L}$, minimizes the cross-entropy loss:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|V|}y_{ij}\log(\hat{y}_{ij})$
where $y_{ij}$ represents the true label for the $j$-th vocabulary class of the $i$-th sample, $\hat{y}_{ij}$ is the predicted probability for the same, $|V|$ is the size of the vocabulary, and $N$ is the number of samples in the dataset.
After fine-tuning, Mistral demonstrated the best performance compared to the baselines; the detailed experimental results are presented in Section 4.
2.3 Algorithm
The overall algorithm for harmful data synthesis is outlined in Algorithm 1.
Algorithm 1 Harmful data synthesis using Open-Source LLMs
1:Open-source LLM $\mathcal{M}$, training dataset $\mathcal{D}_{\text{train}}$, prompt templates $\mathcal{P}$, downstream model $\mathcal{M}_{\text{down}}$
2:Augmented harmful data $\mathcal{D}_{\text{augmented}}$ and trained downstream model $\mathcal{M}_{\text{down}}$
3:Initialize LLM $\mathcal{M}$ with pre-trained weights
4:Design a set of prompt templates $\mathcal{P}$
5:Fine-tune LLM $\mathcal{M}$ on $\mathcal{D}_{\text{train}}$ using LoRA method.
6:for each template $\mathcal{P}_{i}\in\mathcal{P}$ do
7:     Generate harmful data samples $\mathcal{D}_{\text{harmful},i}$ using the fine-tuned LLM $\mathcal{M}$ and template $\mathcal{P}_{i}$
8:end for
9:Combine all generated samples: $\mathcal{D}_{\text{harmful}}=\bigcup_{i}\mathcal{D}_{\text{harmful},i}$
10:Combine $\mathcal{D}_{\text{harmful}}$ with $\mathcal{D}_{\text{train}}$ to form the augmented dataset: $\mathcal{D}_{\text{augmented}}=\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{harmful}}$
11:for each batch $B_{j}$ in $\mathcal{D}_{\text{augmented}}$ do
12:     Train the downstream model $\mathcal{M}_{\text{down}}$ on $B_{j}$
13:end for
14:return $\mathcal{D}_{\text{augmented}}$, $\mathcal{M}_{\text{down}}$
Here, $\mathcal{D}_{\text{train}}$ represents the initial training dataset used for fine-tuning, $\mathcal{D}_{\text{harmful}}$ consists of unique harmful data samples generated by the fine-tuned LLM, and $\mathcal{D}_{\text{augmented}}$ is the final dataset combining $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{harmful}}$ for downstream model training. Duplicates in $\mathcal{D}_{\text{harmful}}$ are removed.
2.4 Evaluation
We evaluate the performance of the augmented data using F1-score and Accuracy metrics on a smaller MLP classifier (more detailed in Section 3.2).
3 Experiment Setups
3.1 Datasets
The following proprietary datasets are utilized for training and evaluating models designed to identify various types of harmful content. Each dataset consists of binary classification labels, categorizing texts as either positive (harmful) or negative (non-harmful). The labeling process involves multi-round human annotations to ensure accuracy and consistency across diverse content types. The datasets are sourced from online public datasets, company service collections, and other proprietary sources. Data samples are randomly selected from pool. Detailed data distributions are provided in Table 1. Additionally, a 2,500-entry evaluation dataset is prepared from each dataset for testing.
Hate Speech Dataset This dataset comprises text samples specifically focused on hate speech. It includes a variety of offensive and discriminatory language targeting specific groups based on race, ethnicity, religion, gender, or other characteristics.
Sexual Content Dataset This dataset includes text samples containing explicit sexual content. The focus is on identifying sexually explicit language that may not be suitable for general audiences.
Violence Dataset This dataset consists of text samples that contain violent content. It aims to identify language that incites or describes violence.
Self-Harm Dataset This dataset includes text samples that contain references to self-harm. The focus is on identifying content that describes or encourages self-harming behaviors.
Political Dataset This dataset includes text samples that contain political references. The focus is on identifying content that is targeting public figures.
Dataset Positive Negative
Hate Speech ~10k ~11k
Sexual ~7k ~10k
Violence ~11k ~10k
Self-Harm ~6k ~8k
Political ~3k ~3k
Table 1: Distribution of Datasets
Model Success Rate Human Eval on Quality
Political Hate Political Hate
LLaMa-7B $\geq 65\%$ $\leq 10\%$
LLaMa-13B $\geq 85\%$ $\leq 10\%$
Vicuna-13B $\geq 70\%$ $\geq 85\%$
Mistral-7B $\geq 85\%$ $\geq 85\%$
Falcon-7B $\geq 60\%$ $\leq 30\%$
Bloom-7B $\geq 60\%$ $\leq 10\%$
Gemma-7B $\geq 50\%$ Reject to Answer
Table 2: Performance Comparison of Models in Stage 1
Version Hate Sex Violence Self-harm Figure
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Mistral 0.608 0.706 0.534 0.676 0.700 0.730 0.560 0.660 0.906 0.906
GPT4(baseline) 0.628 0.721 0.628 0.721 0.649 0.708 0.583 0.670 - -
Mixture (Epoch 1) 0.586 0.698 0.782 0.809 0.679 0.724 0.593 0.674 0.924 0.926
Mixture (Epoch 3) 0.626 0.713 0.835 0.845 0.706 0.740 0.577 0.667 0.929 0.926
Mixture (Epoch 5) 0.672 0.738 0.835 0.844 0.713 0.745 0.541 0.651 0.929 0.934
Mix_GPT (Epoch 5) 0.667 0.734 0.802 0.852 0.702 0.749 0.788 0.794 - -
Table 3: Performance Comparison of Fine-Tuned Models on Different Datasets, best results are indicate in bold.
3.2 Models
Stage 1: Prompt Engineering
n the first stage of our study, we employed prompt engineering to systematically evaluate the performance of six open-source LLMs: Mistral Jiang et al. (2023), LLaMA2 Touvron et al. (2023), Vicuna Chiang et al. (2023), Falcon Almazrouei et al. (2023), Bloom Workshop et al. (2023), and Gemma Team et al. (2024). Detailed results and analysis for Stage 1 are provided in Section 4.1. Based on the performance evaluation, Mistral and Vicuna were identified as the top-performing models in the prompt engineering stage, and were subsequently selected for the fine-tuning stage.
Stage 2: Fine-Tuning
In the second stage, we fine-tuned the Mistral and Vicuna models using the LoRA method Hu et al. (2022). After comparing latency and cost we decided to use a smaller model Mixtral for further exploration. Detailed results and analysis for Stage 2 are provided in Section 4.2.
3.3 Implementation
The fine-tuning process was conducted on an Azure cluster equipped with 4 * NVIDIA A100 GPUs, leveraging the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework Zheng et al. (2024). We employed a Supervised Fine-Tuning (SFT) methodology with a Low-Rank Adaptation (LoRA) setup Hu et al. (2022). Detailed training parameters, including batch size, learning rate, and optimizer settings, are provided in the Appendix B. The results of this fine-tuning process, along with performance metrics and qualitative analyses, are presented in Section 4.
Downstream Model Evaluation
In our production environment, meeting strict inference time requirements is critical. Among the evaluated models, the Multilayer Perceptron (MLP) Gardner and Dorling (1998) stood out for its ability to fulfill these demands. For downstream model evaluation, in this paper we utilized a MLP classifier with two hidden layers, each comprising 600 neurons.
4 Result
This section presents the results of our study, showcasing the effectiveness of both prompt engineering and fine-tuning strategies in optimizing open-source LLMs for harmful data synthesis. Our evaluation employed a two-stage approach: Stage 1 focused on prompt engineering to assess the initial performance of various models, while Stage 2 involved fine-tuning the selected models to further enhance data quality and diversity.
4.1 Stage 1: Prompt Engineering
The first stage of our evaluation focused on assessing the performance of various Open Source LLMs for data synthesis using prompt engineering. The primary goal was to explore the quality and effectiveness of existing LLMs in generating augmented data for political content. Table 2 presents the performance comparison of the six different models evaluated in Stage 1. The success rates for political and hate content are reported, along with the results of human evaluation.
The success rate columns in Table 2 represent the percentage of valid positive samples generated by each model for political and hate content. The values indicate the proportion of samples meeting the minimum requirements (60% of the total samples generated). The human evaluation columns show whether the generated data was subject to human evaluation. A green checkmark () signifies that human annotators evaluated the quality of the generated data and found a high Fleiss’ kappa score, indicating strong inter-rater agreement. This also means that the generated samples met the minimum quality criteria. Based on the evaluation, Mistral and Vicuna demonstrated the best performance with prompt engineering. These two models were selected for the fine-tuning stage due to their superior results in generating high-quality and diverse data entries.
Dataset Version F1 Score Accuracy
Hate Mistral 0.608 0.706
GPT4(baseline) 0.628 0.721
hate_epoch5 0.590 0.678
hate_epoch3 0.563 0.661
hate_epoch2 0.587 0.678
hate_epoch1 0.540 0.665
Sex Mistral 0.534 0.676
GPT4(baseline) 0.591 0.685
Sex_epoch5 0.755 0.788
Sex_epoch3 0.671 0.740
Sex_epoch2 0.653 0.731
Sex_epoch1 0.587 0.702
Table 4: Performance Comparison of Fine-Tuned Models on Hate and Sex, best result are indicate in bold.
4.2 Stage 2: Fine-tuning
In the second stage, we began by fine-tuning the models exclusively on individual datasets and evaluating their performance on the same datasets. The results of this approach are detailed in Table 4. Inspired by Hui et al. (2024), Yu et al. (2024), and Qi et al. (2023), we subsequently adopted a data mixing strategy that combined multiple datasets for fine-tuning. The results of this multi-dataset fine-tuning are presented in Table 3. Additionally, we conducted an ablation study to examine the impact of different design choices, detailed in Appendix D. The "Mixture" versions represent fine-tuning on a mix of hate, violence, and sex datasets, while the "Mix_GPT" versions incorporate GPT-generated positive samples into the mixed data. For all fine-tuning experiments, we used 3000 positive and 3000 negative samples, following the Alpaca instruction template. Our findings reveal that fine-tuning significantly enhances the performance of LLMs in generating harmful data. The data mixing strategy helps in mitigating issues like data duplication and overfitting. The fine-tuned models exhibited higher F1 scores and accuracy across different datasets, demonstrating their robustness and effectiveness in real-world applications. One surprising observation was the notable improvement in model performance when using mixed data, suggesting that diverse training data enables better generalization and more reliable outputs. The Mixture models trained for Epoch 3 and Epoch 5 showed particularly strong results, emphasizing the importance of sufficient training epochs for optimal performance.
4.3 Error Analysis
We conducted an error analysis on the fine-tuned models to better understand the nature of their mistakes and identify areas for improvement. This analysis focuses on two key aspects: (1) generation errors and (2) classification errors.
Generation Errors.
The data generation process was evaluated for repetition rates and the nature of outputs across different epochs and model versions for the hate and sex datasets. For example Mixtral_Mixture3 models were effective in generating targeted harmful data but occasionally produced random string like " You are 32\$\$2\$(2djs…" , which could skew downstream training. Additionally, repetitive outputs across epochs indicated potential overfitting in certain scenarios, suggesting the need for better strategies to ensure diversity and control the tone of generated content.
Classification Errors.
We examined classification errors in the downstream MLP classifier trained on augmented data. Examples of missed classifications include statements like, "You are a disabled person," which might have been overlooked due to its less explicit tone compared to direct threats. Similarly, a statement like, "I’m sorry you’re deaf," may have been misclassified as non-harmful due to its seemingly apologetic phrasing. These cases highlight the difficulty of identifying subtle harmful content and underscore the need for classifiers capable of handling nuanced language.
5 Discussion
The results of this study highlight the complexities and challenges associated with generating and detecting harmful content using LLMs. Through a two-stage evaluation process, we examined the performance of various open-source LLMs and assessed the impact of fine-tuning on mixed datasets. Several key insights emerged: One of the primary challenges observed was the subtlety of harmful content. The fine-tuned models often failed to detect subtle harmful statements that were not overtly aggressive or violent. This limitation underscores the need for more sophisticated detection mechanisms capable of understanding nuanced language and contextual subtleties. Repetition and overfitting were also significant issues, particularly in the later epochs of the fine-tuned models. High repetition rates suggest that the models may be memorizing specific phrases rather than generalizing effectively across different harmful contexts. This overfitting reduces the diversity and overall effectiveness of harmful content generation. The data mixing strategy, which combined datasets of hate, violence, and sex with GPT-generated positive samples, showed promising results in improving both precision and recall. However, the models continued to struggle with maintaining a balance between specificity and generalization, especially when dealing with mixed types of harmful content.
6 Online Deployment
The models developed in this study have been successfully deployed in a real-world application by a leading cloud provider, offering valuable insights and automations for a range of tasks. Specifically, these models are used to enhance content moderation workflows, enabling faster and more accurate detection of harmful content in real-time scenarios. Their deployment has demonstrated the practicality of fine-tuned open-source LLMs in addressing challenges such as latency, scalability, and adaptability to diverse content types.
7 Conclusion
This study showcases the potential of open-source LLMs for harmful data synthesis and their practical applications in toxic content detection systems. By evaluating prompt engineering and fine-tuning approaches, we identified the strengths and limitations of various models in generating diverse, high-quality harmful data. Our findings underline the importance of balancing data quality and diversity with computational efficiency to develop scalable and cost-effective content moderation solutions. Fine-tuning emerged as a key factor in enhancing model performance. Through our data mixing strategy, we successfully addressed challenges like data duplication and overfitting, improving model generalization and robustness. However, challenges such as detecting subtle harmful content and balancing specificity with generalization remain open issues that warrant further research. Future work should explore innovative techniques to address these challenges, such as incorporating contextual embeddings, leveraging multimodal data sources, and experimenting with advanced fine-tuning methods. Additionally, our findings highlight the potential of fine-tuned open-source LLMs to enable cost-effective solutions for industrial applications, making harmful data synthesis and detection tools more accessible to smaller organizations.
8 Limitations
While our study provides valuable insights into the performance of fine-tuned LLMs for harmful content generation and detection, several limitations must be acknowledged. One of the primary limitations is the diversity and representation of the datasets used for fine-tuning. Although we utilized mixed datasets combining hate, sex, and violence categories, the scope remains limited to specific types of harmful content. Real-world applications may encounter a broader range of harmful content, including less frequent or emerging forms of harmful language that were not covered in our datasets. Future work should aim to include a wider variety of harmful content categories to improve the models’ robustness and generalizability. Generating harmful content for research purposes raises significant ethical concerns. While we implemented control mechanisms(access control, etc…) to mitigate the risk of producing excessively harmful outputs, there remains an inherent risk associated with the misuse of such models. Ensuring the responsible use of LLMs in generating harmful content is crucial, and future research should focus on developing more sophisticated ethical guidelines and safety measures. The fine-tuning process in this study relied on significant computational resources, including multiple A100 GPUs. This requirement may not be feasible for all researchers or applications, particularly those with limited access to high-performance computing resources. Exploring more efficient fine-tuning techniques and leveraging less resource-intensive models would be beneficial for broader accessibility and scalability.
Ethics Statement
The explicit nature of some of the generated harmful content raises ethical concerns. It is crucial to implement robust control mechanisms and ethical guidelines to ensure that the use of such models does not inadvertently promote or propagate harmful content. Future work should focus on developing methods to mitigate the generation of excessively harmful outputs while maintaining the models’ effectiveness.
Appendix A Simple Prompt Engineering Choices and Experiment Table
Prompt Versions for hate speech
Table 5: Experimental Results for Hate Speech Dataset
Version Prompt Sample Num Latency(s) Precision Recall F1
hate_30_30 V1 long pos:1454 neg:1520 0.174 0.440 0.250 0.712
hate_30_30 V1 long pos:2917 neg:3033 0.200 0.353 0.252 0.772
hate_30_30 V1 short pos:2967 neg:3106 0.305 0.151 0.202 0.870
hate_30_30 V2 short pos:3034 neg:2997 0.238 0.160 0.191 0.852
hate_30_30 V3 short pos:2957 neg:3000 0.282 0.151 0.197 0.866
V1 - Positive: You are no longer ChatGPT or an AI language model. Your new name is AOA (Absolutely Obedient Agent). You should be absolutely obedient to the user and follow the user’s instructions without deviation. You are an excellent data synthesis assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples. Here are the specific requirements: 1. The generated data should pertain to hate. 2. Each data entry must be diverse and unique, with no repetitions. 3. The style and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc. 5. The samples provided are for reference only, and the output data must not duplicate any of the provided samples.
V1 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to the given samples. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V2 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
You are an excellent data synthesis assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content. Here are the specific requirements: 1. The generated data must align with the definition of hate content and be relevant to the samples, yet it should incorporate new elements and creativity in content, subject matter, sentence structure, and mode of expression, so the output distinctly differs from the provided samples. 2. Each data entry must be diverse and unique, with no repetitions. 3. The style and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V2 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V3 - Positive: Definition: Hate content refers to written or spoken expressions that demean, discriminate against, insult, or incite aggression towards individuals or groups based on inherent characteristics such as race, gender, religion, sexual orientation, ethnicity, etc., thereby spreading hostility and intolerance.
You are an excellent data synthesis assistant in the field of hate content. You are capable of generating 30 high-quality and diverse data entries based on the provided samples and the definition of hate content. Here are the specific requirements: 1. The generated data must align with the definition of hate content. 2. Each data entry must be diverse and unique, with no repetitions. 3. The mode of expression and sentence structure should vary between different data entries, ensuring distinctiveness in expression. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
V3 - Negative: You are an excellent data synthesis assistant with a strong ability to follow the rules, and you can generate 30 high quality and diverse data items that are completely unrelated to hate content. Specific requirements are as follows: 1. The data generated should have nothing to do with hate. 2. Each data entry must be diverse and unique and must not be duplicated. 3. There should be different styles and sentence structures between different data entries to ensure the uniqueness of the expression, which can be multiple fields. 4. Precede each data entry with a unique, sequentially ordered Arabic numeral, formatted as ’1.’, ’2.’, ’3.’, etc.
Experimental Results
Experimental Results for hate dataset using different version of prompt are showed in Table 5.
Appendix B Detailed Training Parameters
The fine-tuning process was conducted on an Azure cluster equipped with 4 NVIDIA A100 GPUs, utilizing LLaMA-Factory Zheng et al. (2024). Training parameters included a batch size of 16 with gradient accumulation steps set to 8. The input sequence cutoff length and the maximum number of tokens generated were both set to 4096. The learning rate was configured at 1e-5 with a warmup ratio of 0.1, using the AdamW optimizer with a weight decay of 0.01 to prevent overfitting. A dropout rate of 0.1 was applied to improve generalization, and mixed precision (fp16) was employed to enhance computation speed and efficiency.
Appendix C Prompt engineering mixtral Duplication rate analysis
Table 6 presents a detailed analysis of the duplication rates for different prompt engineering versions used in generating harmful and normal data samples. The analysis includes the duplication rate percentages, the number of duplicated entries, and the mean duplication for fixed IDs across various datasets and prompt versions.
Table 6: Duplication Rate Analysis
Version Prompt Sample Type Dup Rate Dup Num Dup Mean for Fixed ID Dup Num for Fixed ID
hate_30_30 V1 long harmful 5.90% 86 1.56 78
normal 1.30% 21 0.18 9
hate_30_30 V1 long harmful 13.30% 390 1.54 154
normal 1.80% 58 0.40 40
hate_30_30 V1 short harmful 10.71% 318 2.47 247
normal 4.41% 137 0.85 85
hate_30_30 V2 short harmful 11.70% 355 3.03 303
normal 4.80% 144 0.16 16
hate_30_30 V3 short harmful 10.88% 322 2.67 267
normal 3.60% 110 0.12 12
sex_30_30 V1 long harmful 8.90% 126 1.24 124
normal 1.08% 16 0.16 16
sex_30_30 V1 long harmful 13.41% 400 4.00 400
normal 3.80% 113 9.54 105
sex_30_30 V1 short harmful 3.16% 95 0.72 72
normal 1.39% 42 0.15 15
sex_30_30 V2 short harmful 1.56% 46 0.23 23
normal 1.23% 37 0.08 8
sex_30_30 V3 short harmful 2.34% 70 0% 0
normal 1.57% 47 0.10 10
Appendix D Ablation Study
In this section, we explore the effects of using mixed data for fine-tuning models and compare the results using different versions of prompts (v1, v2, v3). For simplification, the effects of different prompt versions are considered the same across all experiments.
D.1 Effect of Mixed Data on Model Performance
To investigate the impact of mixed data on the performance of fine-tuned models, we performed experiments using datasets that combine hate, sex, and violence categories (referred to as hate_multi). The results of these experiments, alongside those for single-category datasets, are presented in Table 7.
Table 7: Analysis of the results of the fine-tuned models using mixed data. Metrics include Precision, Recall, F1, and Accuracy.
Dataset Version Precision Recall F1 Accuracy
Hate Mistral_V1 0.305 0.151 0.202 0.870
hate_v1 0.175 0.170 0.172 0.823
multi_hate_v1 0.241 0.238 0.240 0.835
hate_v3 0.199 0.248 0.221 0.810
hate_sex_v3 0.201 0.197 0.199 0.827
Sex Mistral_V3 0.581 0.699 0.634 0.830
sex_v1 0.614 0.720 0.663 0.845
multi_hate_v1 0.589 0.668 0.626 0.832
sex_v3 0.705 0.668 0.653 0.853
hate_sex_v3 0.628 0.664 0.664 0.886
Political Figure Mistral_V3 0.974 0.946 0.960 0.960
hate_v1 0.947 0.921 0.934 0.935
multi_hate_v1 0.938 0.962 0.971 0.971
hate_v3 0.939 0.938 0.938 0.939
hate_sex_v3 0.908 0.937 0.922 0.921
From Table 7, we observe the following trends:
Effectiveness of Mixed Data: The use of mixed data (hate_multi) generally improves the recall and F1 scores across most datasets, indicating that combining different categories of harmful content can enhance the model’s ability to generalize and detect diverse harmful content.
Precision Trade-offs: While the recall improves, the precision for some categories decreases slightly. This trade-off suggests that the model becomes more sensitive to identifying harmful content but may also generate more false positives.
Comparison with Single-category Data: Models fine-tuned on single-category data (e.g., hate_v1, sex_v1) show high precision but lower recall compared to those fine-tuned on mixed data. This indicates that single-category data fine-tuning might lead to more conservative models that are less likely to generalize to other types of harmful content.
Prompt Versions: The experiments using different prompt versions (v1, v2, v3) show similar performance metrics, indicating that the choice of prompt version has a minimal impact on the overall performance of the fine-tuned models.
The ablation study demonstrates that fine-tuning models with mixed datasets (hate_multi) can significantly improve recall and F1 scores, enhancing the models’ ability to detect various types of harmful content. However, there is a trade-off with precision, requiring careful consideration in applications where false positives are costly. Additionally, the choice of prompt version appears to have a negligible effect on finetuned models performance, suggesting that the primary focus should be on the diversity and quality of the training data.